Ricardo has a monthly budget of $20 to spend on movie tickets and wine. The price of a movie tickets is $10 and the price of a bottle of wine is $5. a. Calculate the relative price of wine in terms of movie tickets? (1 mark)

b. Calculate Ricardo’s real income in terms of wine? (2 mark)

c. Calculate Ricardo’s real income in terms of movie tickets? (2 mark)


d. Calculate the equation of Ricardos budget line with wine on the left side. (3marks).


e. Draw a graph for Ricardos budget line with quantity of movie tickets on the X-axis and quantity of wine on the Y axis. (2 marks).
1
Expert's answer
2019-06-10T09:09:20-0400

a. The relative price of wine in terms of movie tickets is 5/10 = 0.5 movie ticket.

b. Ricardo’s real income in terms of wine is 20/5 = 4 bottles.

c. Ricardo’s real income in terms of movie tickets is 20/10 = 2 movie tickets.

d. The equation of Ricardos budget line with wine on the left side is I = 5Qw + 10Qm = 20.

e. A graph for Ricardos budget line with quantity of movie tickets on the X-axis and quantity of wine on the Y axis is a representation of the equation from d.